wht would be percentage composition by volume of a mixture of CO and CH4 whose 10.5 ml requires 9 ml oxygen for combution?


1)CO=80% CH4=20% 2)CO =90% CH4 =10%


3) CO=76.2% CH4=23.8% 4) CO=66% CH4=34%

Grade:12

1 Answers

Aman Bansal
11 years ago

Dear Soumya,

Moles = pV / RT = 1.00 atm x 1.00 L / 0.0821 x 273.15 = 0.0446

let x = moles CH4 and let y = moles C2H6

x + y = 0.0446
x ( 890.3) + y ( 1559.7) = 46.9

x = 0.0446 -y
39.7 - 890.3 y + 1559.7 y = 46.9

669.4 y = 7.2

y = 0.0108 moles C2H6
x = 0.0446 - 0.0108 = 0.0338

V(C2H6) = 0.0338 x 0.0821 x 273.15 / 1.00 =0.758 L
V ( CH4) = 0.0108 x 0.0821 x 273.15 / 1.00 =0.242 L

% by volume = 0.242 x 100 / 1.00 = 24.2 ( CH4)
and 75.8 % C2H6

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert